OpenStudy (anonymous):

Solve for indicated variable in this equation 3x-2y=18 solve for x

OpenStudy (anonymous):

First move -2y to the other side of the equation, the sign will flip and it becomes: 3x = 18 + 2y divide both sides by 3: x = (18 + 2y)/3

OpenStudy (anonymous):

Does that makes sense?